# Critical Quantum Sensing: a tutorial on parameter estimation near quantum phase transitions

### v1

@@ IX Driven-dissipative Kerr resonator
Accordingly, we will show that critical protocols based on the system’s steady state can achieve at most a linear scaling of the QFI. However, we will then show how the dynamical critical properties can be exploited to achieve the Heisenberg scaling for finite evolution times. In particular, in Sec. IX.4, we analyze the transition between the Heisenberg and the dissipative regimes. Finally, in Sec. IX.5, we show how to identify the optimal working regime of a dissipative critical protocol for finite resources.
@@
where $\omega_{r}$ is the resonance frequency, $\chi$ is the Kerr nonlinearity and $\epsilon$ is the intensity of the two-photon pump of frequency $\omega_{p}$. Nonlinear quantum-optical oscillators represented a key testbed for the study of finite-component first-order dissipative phase transitions, both from the experimental [230, 231, 232, 233] and the theoretical viewpoints [234, 235, 236, 237]. Similar parametric systems have also been considered in sensing applications, such as optical parametric oscillators  [238, 239, 240]. Here, we consider in particular the case of parametric (two-photon) pump, where a second-order dissipative phase transition has recently been observed [226, 65].
@@ IX.2 Static protocol
Let us first analyze the metrological performance in terms of the scaling, with respect to fundamental resources, of the QFI calculated over the steady-state manifold. As shown in the previous subsection, the steady state of the considered model is Gaussian. We can then use a general formula for the QFI that holds for arbitrary single-mode Gaussian state manifolds. For the estimation of a single parameter $\omega$, such formula can be computed from first-moments vector $\bf v$ and covariance matrix $\bf\Sigma$} [242],
@@
where $\mu=1/\sqrt{\det\left({\bf\Sigma}\right)}$ is the purity of the Gaussian state. We focus on the estimation of the parameter $\omega$ in Eq. (197), assuming that all the other Hamiltonian and dissipative parameters are known. We consider the case of local parameter estimation—as defined in Sec. (III)—and so we assume that $\omega=\omega_{0}+\delta\omega$, where $\omega_{0}$ is the prior information over the parameter and $\delta\omega$ is a small frequency shift we aim to estimate. In this case, the QFI in Eq. (206), is evaluated by taking the derivatives and then evaluating the resulting function at $\omega=\omega_{0}$. Accordingly, we derive the QFI for the estimation of $\omega$ by substituting the steady-state solutions of Eq. (204) into Eq. (206), yielding
@@
Notice that, if $\omega_{0}\neq 0$, when $\epsilon\to\epsilon_{c}$ the QFI achieves a quadratic scaling with respect to the number of photons, i.e., $\mathcal{I}_{\omega}^{ss}\sim 8\omega_{0}^{2}N^{2}_{ss}/\epsilon^{4}$. However, this quadratic scaling does not take into account the time $T_{ss}$ required to reach the steady state. As discussed in Sec.IX.3, it can be lower bounded by the inverse of the smallest eigenvalue of the dynamics matrix $T_{ss}\gtrsim\lambda_{-}^{-1}$. We have that $\lambda_{-}^{-1}\sim\Gamma/\epsilon_{c}(\epsilon_{c}-\epsilon)$, in the limit $\epsilon\to\epsilon_{c}$. We then find
@@ IX.3 Gaussian approximation: dynamics
So far, our discussion has been restricted to steady-state properties, where the system has already relaxed under the combined action of drive and dissipation. However, it is equally important—and often more practically relevant—to ask how the metrological performance evolves during the transient dynamics, before the system settles into its steady state. In what follows, we explore this time-dependent behavior and assess the potential of dynamical protocols for critical quantum sensing. Under Gaussian approximation, analytical solutions for the system’s dynamics can be found by solving the Langevin equations [243]:
@@
Equation (211) describes the evolution of the resonator field $\hat{a}$, where the quantum noise of the Markovian environment is described by the input mode $\hat{\textbf{b}}(t)=(\hat{b}(t),\hat{b}^{\dagger}(t))^{\rm T}$. Since we are working at zero temperature, and we have already included the parametric driving as an Hamiltonian term, we assume the input mode to be the vacuum, such that $\langle\hat{b}^{\dagger}(t)\hat{b}(t^{\prime})\rangle=\delta(t^{\prime}-t)$. Exact solutions to the linear system of first-order differential equations (211) can be written as [59]:
@@
where $\lambda_{\pm}=\Gamma\pm\sqrt{\epsilon^{2}-\omega^{2}}$ are the eigenvalues of the dynamics matrix A-A. To derive the time-dependent covariance matrix of the system, ${\bf{\Sigma}}(t)$, we can express the position and momentum operators as functions of time using Eq. (213). We emphasize that solution from Eq. (213) is valid only below the critical point, i.e., when $\epsilon<\epsilon_{c}$. Under Gaussian approximation, analytical time-dependent solutions can also be found in the regime $\epsilon>\epsilon_{c}$. Analogously to the treatment discussed in Sec. IX.1, a Gaussian description beyond the critical point can be found by considering quantum fluctuations around semiclassical solutions. Since our main interest lies in understanding the behavior as the critical point is approached, in the following we will focus on the case $\epsilon<\epsilon_{c}$.
Figure 15: Optimal measurement strategy of dynamical protocols. The optimal sensitivity QFI is shown in the blue line and compared with the Fisher information of the homodyne detection optimized over the quadrature angle shown by the red dots. (a) Shows the unitary regime with $\epsilon=0.99\omega_{0}$ and (b) shows the dissipative regime with $\omega_{0}=\Gamma,\,N_{\rm max}=100,\,\epsilon=\epsilon_{\rm opt}=\sqrt{2N_{\rm max}/(1+2N_{\rm max})}\epsilon_{c}$. The optimized homodyne detection is shown to saturate the QFI.
Before proceeding, let us briefly examine the dynamics of the photon number inside the cavity, assuming the system is initialized in the vacuum. The average number of photons at time $t$ is given by $N(t)=\langle\hat{a}^{\dagger}(t)\hat{a}(t)\rangle$ where $\hat{a}(t)$ is defined in Eq. (213). Depending on the value of the pump strength $\epsilon$, two distinct behaviors of $N(t)$ emerge. For $\epsilon<\omega$, the average photon number is given by
@@ IX.4 Dynamical protocol: from Heisenberg to dissipative regime
Having established an analytical description of the system’s time-dependent dynamics, we are now in a position to return to our central task: parameter estimation. The transient evolution not only enriches our understanding of critical behavior but also provides a natural stage to investigate how information about the parameters of interest is encoded in the system at different times. In the following, we shift our focus back to estimation theory and analyze how these dynamical features can be harnessed for quantum-enhanced sensing. The time-dependent solutions given in Eq. (213) allow us to compute the instantaneous QFI $\mathcal{I}_{\omega}^{t}$. The latter has to be interpreted as the QFI of a protocol where the system is evolved until time $t$ and then measured. In particular, by substituting Eq. (213) into Eq. (202) and using the general expression for the QFI of single-mode Gaussian states of Eq. (206), we can directly evaluate $\mathcal{I}_{\omega}^{t}$. Although its full expression is too lengthy to report here, in the following we will analyze how it scales with the available resources in different regimes of interest.
We begin by examining the simpler case where we set $\Gamma=0$, and the resonator evolves following a unitary dynamics. The critical point simplifies to $\epsilon_{c}=\omega$. The covariance matrix of the system in this regime is given by
@@ IX.5 Optimization of dynamical protocol in the dissipative regime
Figure 16: Instantaneous QFI of the dynamical protocols. (a) Shows the time-dependent QFI of the dynamical protocol $\mathcal{I}_{\omega}^{t}$. (b) Shows the ratio $\mathcal{I}_{\omega}^{t}/N^{2}(t)$ as a function of time. In the presence of dissipation, the single-shot QFI scales quadratically with the number of photons—achieving Heisenberg scaling. This is attributed to the emergence of a transient regime as $\Gamma>0$, during which the QFI can increase. The parameters are chosen such that $\omega_{0}=\Gamma,\,N_{\rm max}=100,\epsilon=\epsilon_{\rm opt}=\sqrt{2N_{\rm max}/(1+2N_{\rm max})}\epsilon_{c}$.
@@ X Critical metrology in ultrastrongly coupled light-matter systems
In the previous section, we examined a driven-dissipative mechanism in which an external drive generates squeezing in the system’s steady state. We now shift focus to a fundamentally different process—interaction-induced frequency renormalization—where the internal interactions among the system’s constituents effectively modify, or squeeze, its natural frequency [244]. This effect is particularly prominent in cavity quantum electrodynamics, where embedding atoms or other media inside an optical cavity leads to a shift in the cavity’s resonance frequency due to light-matter coupling.
@@ X.4 Finite-size effects
The calculations in this section assume the limit $\omega/\Omega\rightarrow 0$, where the system reduces to a harmonic oscillator with a tunable frequency. This captures the essential physics and predicts maximal squeezing, but it is an idealization. For any finite ratio $\omega/\Omega$, anharmonicities appear, leading to nonlinear effects such as photon blockade [274], which limit the achievable coherent field amplitude. Moreover, realistic systems are finite, so the energy gap never fully closes, bounding both the QFI and the accessible squeezing parameter $\xi$. Thus, the divergence of $\xi$ predicted in the strict limit cannot be realized in practice. A way around these limitations is to move to the many-body regime described by the Dicke model [275], where a single bosonic mode couples to $N$ two-level atoms. For large $N$, the collective spin acts almost classically, restoring effectively harmonic dynamics near resonance without requiring extreme detuning. This collective enhancement provides a practical route to strong squeezing and metrological advantages that are inaccessible in single-atom setups.
@@
XI Conclusions

### v2

@@ IX Driven-dissipative Kerr resonator
Accordingly, we will show that critical protocols based on the system’s steady state can achieve at most a linear scaling of the QFI. However, we will then show how the dynamical critical properties can be exploited to achieve the Heisenberg scaling for finite evolution times. In particular, in Section IX.4, we analyse the transition between the Heisenberg and the dissipative regimes. Finally, in Section IX.5, we show how to identify the optimal working regime of a dissipative critical protocol for finite resources.
@@
where $\omega_{r}$ is the resonance frequency, $\chi$ is the Kerr nonlinearity, and $\epsilon$ is the intensity of the two-photon pump of frequency $\omega_{p}$. Nonlinear quantum-optical oscillators represented a key testbed for the study of finite-component first-order dissipative phase transitions, both from the experimental [245, 246, 247, 248] and the theoretical viewpoints [249, 250, 251, 252]. Similar parametric systems have also been considered in sensing applications, such as optical parametric oscillators  [253, 254, 255]. Here, we consider in particular the case of parametric (two-photon) pump, where a second-order dissipative phase transition has recently been observed [241, 68].
@@ IX.2 Static protocol
Let us first analyze the metrological performance in terms of the scaling, with respect to fundamental resources, of the QFI calculated over the steady-state manifold. As shown in the previous subsection, the steady state of the considered model is Gaussian. We can then use a general formula for the QFI that holds for arbitrary single-mode Gaussian state manifolds. For the estimation of a single parameter $\omega$, such formula can be computed from first-moments vector $\bf v$ and covariance matrix $\bf\Sigma$} [257]:
@@
where $\mu=1/\sqrt{\det\left({\bf\Sigma}\right)}$ is the purity of the Gaussian state. We focus on the estimation of the parameter $\omega$ in Eq. (197), assuming that all the other Hamiltonian and dissipative parameters are known. We consider the case of local parameter estimation—as defined in Section (III)—and so we assume that $\omega=\omega_{0}+\delta\omega$, where $\omega_{0}$ is the prior information over the parameter and $\delta\omega$ is a small frequency shift we aim to estimate. In this case, the QFI in Eq. (206) is evaluated by taking the derivatives and then evaluating the resulting function at $\omega=\omega_{0}$. Accordingly, we derive the QFI for the estimation of $\omega$ by substituting the steady-state solutions of Eq. (204) into Eq. (206), yielding
@@
Notice that, if $\omega_{0}\neq 0$, when $\epsilon\to\epsilon_{c}$ the QFI achieves a quadratic scaling with respect to the number of photons, i.e., $\mathcal{I}_{\omega}^{ss}\sim 8\omega_{0}^{2}N^{2}_{ss}/\epsilon^{4}$. However, this quadratic scaling does not take into account the time $T_{ss}$ required to reach the steady state. As discussed in Section IX.3, it can be lower bounded by the inverse of the smallest eigenvalue of the dynamics matrix $T_{ss}\gtrsim\lambda_{-}^{-1}$. We have that $\lambda_{-}^{-1}\sim\Gamma/\epsilon_{c}(\epsilon_{c}-\epsilon)$, in the limit $\epsilon\to\epsilon_{c}$. We then find
@@ IX.3 Gaussian approximation: dynamics
So far, our discussion has been restricted to steady-state properties, where the system has already relaxed under the combined action of drive and dissipation. However, it is equally important—and often more practically relevant—to ask how the metrological performance evolves during the transient dynamics, before the system settles into its steady state. In what follows, we explore this time-dependent behaviour and assess the potential of dynamical protocols for critical quantum sensing. Under Gaussian approximation, analytical solutions for the system’s dynamics can be found by solving the Langevin equations [258]:
@@
Equation (211) describes the evolution of the resonator field $\hat{a}$, where the quantum noise of the Markovian environment is described by the input mode $\hat{\textbf{b}}(t)=(\hat{b}(t),\hat{b}^{\dagger}(t))^{\rm T}$. Since we are working at zero temperature and we have already included the parametric driving as an Hamiltonian term, we assume the input mode to be the vacuum, such that $\langle\hat{b}^{\dagger}(t)\hat{b}(t^{\prime})\rangle=\delta(t^{\prime}-t)$. Exact solutions to the linear system of first-order differential equations (211) can be written as [62]:
@@
where $\lambda_{\pm}=\Gamma\pm\sqrt{\epsilon^{2}-\omega^{2}}$ are the eigenvalues of the dynamics matrix A-A. To derive the time-dependent covariance matrix of the system, ${\bf{\Sigma}}(t)$, we can express the position and momentum operators as functions of time using Eq. (213). We emphasize that solution from Eq. (213) is valid only below the critical point, i.e., when $\epsilon<\epsilon_{c}$. Under Gaussian approximation, analytical time-dependent solutions can also be found in the regime $\epsilon>\epsilon_{c}$. Analogously to the treatment discussed in Section IX.1, a Gaussian description beyond the critical point can be found by considering quantum fluctuations around semiclassical solutions. Since our main interest lies in understanding the behavior as the critical point is approached, in the following we will focus on the case $\epsilon<\epsilon_{c}$.
Figure 15: Optimal measurement strategy of dynamical protocols. We compare the QFI (blue line) and the Fisher information of the homodyne detection (red dots), optimized over the quadrature angle. Panel (a) shows the unitary regime ($\Gamma=0$) with $\epsilon=0.99\omega_{0}$, while (b) the dissipative regime with $\omega_{0}=\Gamma,\,N_{\rm max}=100,\,\epsilon=\epsilon_{\rm opt}=\sqrt{2N_{\rm max}/(1+2N_{\rm max})}\epsilon_{c}$. The optimized homodyne detection is shown to saturate the QFI.
Before proceeding, let us briefly examine the dynamics of the photon number inside the cavity, assuming the system is initialized in the vacuum. The average number of photons at time $t$ is given by $N(t)=\langle\hat{a}^{\dagger}(t)\hat{a}(t)\rangle$ where $\hat{a}(t)$ is defined in Eq. (213). Depending on the value of the pump strength $\epsilon$, two distinct behaviours of $N(t)$ emerge. For $\epsilon<\omega$, the average photon number is given by
@@ IX.4 Dynamical protocol: from Heisenberg to dissipative regime
Having established an analytical description of the system’s time-dependent dynamics, we are now in a position to return to our central task: parameter estimation. The transient evolution not only enriches our understanding of critical behavior but also provides a natural stage to investigate how information about the parameters of interest is encoded in the system at different times. In the following, we shift our focus back to estimation theory and analyze how these dynamical features can be harnessed for quantum-enhanced sensing. The time-dependent solutions given in Eq. (213) allow us to compute the instantaneous QFI, $\mathcal{I}_{\omega}^{t}$. The latter has to be interpreted as the QFI of a protocol where the system is evolved until time $t$ and then measured. In particular, by substituting Eq. (213) into Eq. (202) and using the general expression for the QFI of single-mode Gaussian states of Eq. (206), we can directly evaluate $\mathcal{I}_{\omega}^{t}$. Although its full expression is too lengthy to report here, in the following we will analyse how it scales with the available resources in different regimes of interest.
We begin by examining the simpler case where we set $\Gamma=0$, and the resonator evolves following a unitary dynamics. Consequently, the critical point simplifies to $\epsilon_{c}=\omega$. The covariance matrix of the system in this regime is given by
@@ IX.5 Optimization of dynamical protocol in the dissipative regime
Figure 16: Instantaneous QFI of the dynamical protocol. Panel (a) shows the time-dependent QFI of the dynamical protocol, $\mathcal{I}_{\omega}^{t}$, while (b) shows the ratio $\mathcal{I}_{\omega}^{t}/N^{2}(t)$, as a function of time. In the presence of dissipation, the single-shot QFI scales quadratically with the number of photons—achieving Heisenberg scaling. This is attributed to the emergence of a transient regime as $\Gamma>0$, during which the QFI can increase. The parameters are chosen such that $\omega_{0}=\Gamma,\,N_{\rm max}=100,\epsilon=\epsilon_{\rm opt}=\sqrt{2N_{\rm max}/(1+2N_{\rm max})}\epsilon_{c}$.
@@ X Critical metrology in ultrastrongly coupled light-matter systems
In the previous section, we examined a driven-dissipative mechanism in which an external drive generates squeezing in the system’s steady state. We now shift focus to a fundamentally different process: interaction-induced frequency renormalization, where the internal interactions among the system’s constituents effectively modify, or squeeze, its natural frequency [259]. This effect is particularly prominent in cavity quantum electrodynamics, where embedding atoms or other media inside an optical cavity leads to a shift in the cavity’s resonance frequency due to light-matter coupling.
@@ X.4 Finite-size effects
The calculations above assume the limit $\omega/\Omega\rightarrow 0$, where the system reduces to a harmonic oscillator with a tunable frequency. This captures the essential physics and predicts maximal squeezing, but it is an idealization. For any finite ratio $\Omega/\omega$, anharmonicities appear, leading to nonlinear effects such as photon blockade [289], which limit the achievable coherent field amplitude. Moreover, realistic systems are finite, so the energy gap never fully closes, bounding both the QFI and the accessible squeezing parameter $\xi$. Thus, the divergence of $\xi$ predicted in the strict limit cannot be realized in practice. A way around these limitations is to move to the many-body regime described by the Dicke model [290], where a single bosonic mode couples to $N$ two-level atoms. For large $N$, the collective spin acts almost classically, restoring effectively harmonic dynamics near resonance without requiring extreme detuning. This collective enhancement provides a practical route to strong squeezing and metrological advantages that are inaccessible in single-atom setups.
@@
XI Critical metrology in strongly correlated electron systems
As a final example, we now turn to a fundamentally different platform for critical quantum sensing: fermionic impurity systems hosting boundary quantum critical points. These systems constitute paradigmatic examples of strongly non-Markovian open quantum systems, where extensive multipartite entanglement between the impurity and its reservoirs builds up at low temperatures. As a result, perturbative treatments—such as those applicable to the driven dissipative Kerr resonator and ultrastrong light-matter platforms discussed previously—fail to capture the essential correlations and critical behavior. Remarkably, however, these non-Markovian open-system effects can be treated exactly: either numerically by employing the numerical renormalization group (NRG) originally introduced by Wilson [174], or analytically at special solvable points of the model [291, 292]. To illustrate these ideas concretely, we will consider the exactly solvable point of the two-channel Kondo model (2CK), where the strongly correlated electron physics at the critical point can be mapped onto a non-interacting theory of free fermions, reducing a typically difficult problem into a pedagogical solvable model.
XI.1 Two-channel Kondo model
The Kondo model is one of the paradigmatic models of strongly correlated electron physics [293]. Originally introduced to explain the anomalous increase in resistivity observed in metals containing dilute magnetic impurities, it captures how a localized magnetic moment interacts with a sea of conduction electrons via antiferromagnetic exchange. Below an emergent low-energy scale, known as the Kondo temperature $T_{K}$, the impurity spin becomes dynamically screened by the surrounding electrons, forming a non-trivial many-body singlet state. This correlated screening process accounts for the characteristic rise in resistivity at low temperatures arising from enhanced spin-flip scattering of conduction electrons off the impurity.
From a modern perspective, Kondo physics has found a renewed relevance in nanoscale electronic systems. Semiconductor quantum dots, single-molecule transistors, and hybrid superconductor-semiconductor devices can all realize controllable Kondo effects, where a localized electronic level with strong Coulomb interaction acts as an effective magnetic impurity coupled to metallic leads [294, 295, 295]. In such settings, parameters such as gate voltage, coupling asymmetry, and local magnetic fields can be tuned in situ, allowing detailed experimental exploration of strongly correlated phenomena originally discovered in bulk systems. These include the emergence of SU(4) and orbital Kondo effects, spin-orbit-entangled screening, and even Majorana-Kondo hybridization in nanowires [296, 297, 298, 299]. Importantly, these effects do not rely on delicate fine-tuning, but rather, emerge inherently from strong electron-electron interactions, reflecting universal many-body phenomena which appear across a wide range of materials and devices. The resulting collective states host robust quantum coherence, entanglement, and enhanced response functions—resources of direct relevance to quantum metrology and quantum information processing, where sensitivity near correlated critical points can be fundamentally amplified. The Kondo model thus continues to serve as both a conceptual test-bed for the study of strong correlations and a practical platform for realizing correlated quantum technologies in mesoscopic devices.
The two-channel Kondo (2CK) model represents one of the most celebrated examples of non-Fermi liquid physics and boundary quantum criticality in strongly correlated electron systems [300]. Originally introduced as an extension of the conventional Kondo problem, it describes a localized spin-$\tfrac{1}{2}$ impurity symmetrically coupled to two independent conduction electron channels (or “leads”) whose Hamiltonian reads
$\hat{H}_{2CK}=\hat{H}_{\text{leads}}+J_{L}\hat{\mathbf{S}}\cdot\hat{\mathbf{s}}_{0L}+J_{R}\hat{\mathbf{S}}\cdot\hat{\mathbf{s}}_{0R}\;,$ (237)
with $\hat{\mathbf{S}}$ a spin-$\tfrac{1}{2}$ operator for a single impurity, exchange coupled to two independent conduction electron channels with antiferromagnetic coupling $J_{\alpha}>0$. The impurity couples locally to the spin density of conduction electrons in each channel at the impurity site $\hat{\mathbf{s}}_{0\alpha}=\tfrac{1}{2}\sum_{kk^{\prime}\sigma\sigma^{\prime}}\hat{c}^{\dagger}_{\alpha\sigma k}\boldsymbol{\sigma}_{\sigma\sigma^{\prime}}^{\phantom{\dagger}}\hat{c}^{\phantom{\dagger}}_{\alpha\sigma^{\prime}k^{\prime}}$ where $\boldsymbol{\sigma}_{\sigma\sigma^{\prime}}$ is a matrix of Pauli matrices. The leads are given by
$\hat{H}_{\text{leads}}=\sum_{\alpha\sigma k}\epsilon_{k}^{\phantom{\dagger}}\hat{c}^{\dagger}_{\alpha\sigma k}\hat{c}^{\phantom{\dagger}}_{\alpha\sigma k}\;,$ (238)
[…]
Experimentally, realising the two-channel Kondo effect poses significant challenges. This is a consequence of the difficulty in engineering two independent symmetrically coupled electronic leads that compete to screen a single localized spin. Nonetheless, remarkable progress has been achieved in mesoscopic devices that emulate such boundary criticality. Semiconductor quantum dots coupled to metallic leads have demonstrated signatures consistent with overscreened behaviour: such as non-Fermi liquid scaling of conductance and residual zero-bias anomalies [301]. At the same time, the theoretical 2CK model admits an exactly solvable point—corresponding to a special symmetry limit in which the strongly correlated problem can be mapped onto a non-interacting theory of free fermions—providing a rare analytic window into the quantum critical behavior that these experiments seek to emulate [291]. We will turn to this solvable limit in the following section to illustrate explicitly how non-Fermi-liquid correlations and boundary criticality manifest in an exactly tractable form.
XI.2 Exactly solvable point of the model
The intrinsic difficulty of the 2CK problem originates from the presence of two equivalent screening channels competing to screen a single spin-$\tfrac{1}{2}$ impurity. In the frustrated, symmetric coupling regime $J_{L}=J_{R}$, extensive entanglement between the impurity and a macroscopic number of fermionic degrees of freedom in the leads gives rise to a non-Fermi liquid ground state which cannot be accessed perturbatively in either the weak or strong-coupling limits. Remarkably, however, a special limit of the 2CK model—the so called Emery-Kivelson point (EK)—admits an exact analytical solution. This point corresponds to a fine-tuned anisotropic limit of the exchange couplings $J_{\alpha}$ in which the transverse and longitudinal spin-exchange components satisfy a particular relation allowing an exact mapping of the strongly interacting Hamiltonian onto a quadratic model of fermions.
The standard approach to this mapping employs bosonization of the conduction-electron fields, followed by a canonical transformation that diagonalizes the longitudinal exchange term. After refermionization of the remaining bosonic modes, the system reduces to an effective Hamiltonian in which only one linear combination of electronic fields couples to the impurity spin. At the EK point, the spin-flip term becomes bilinear in these new fermionic degrees of freedom, and the resulting model can be expressed as a Majorana resonant-level model (MRLM) describing a single localized Majorana mode hybridized with a continuum of conduction-electron Majoranas. In its simplest form, the effective low-energy Hamiltonian at the EK point reads [291]
$\hat{H}_{\text{MRLM}}=\epsilon_{d}\hat{d}^{\dagger}\hat{d}^{\phantom{\dagger}}+\sum_{k}\epsilon_{k}\hat{c}_{k}^{\dagger}\hat{c}^{\phantom{\dagger}}_{k}+V\sum_{k}\big(\hat{c}^{\dagger}_{k}+\hat{c}_{k}^{\phantom{\dagger}}\big)\big(\hat{d}^{\dagger}-\hat{d}^{\phantom{\dagger}}\big)\;,$ (239)
where $\hat{c}_{k}$ ($\hat{c}^{\dagger}_{k}$ ) annihilates (creates) a conduction-electron with energy $\epsilon_{k}$, $\hat{d}$ describes a localized impurity fermion of energy $\epsilon_{d}$ and $V$ is the hybridization. The key feature of this Hamiltonian is the non-standard hybridization which couples specific Majorana components of the conduction and impurity fermions, rather than their full complex forms. By defining the Majorana operators,
$\hat{\eta}_{k}=\frac{1}{\sqrt{2}}\big(\hat{c}_{k}^{\phantom{\dagger}}+\hat{c}_{k}^{\dagger}\big)\;,\qquad\hat{\xi}_{k}=\frac{i}{\sqrt{2}}\big(\hat{c}_{k}^{\dagger}-\hat{c}_{k}\big)\;,$ (240)
and similarly,
[…]
one can obtain Majorana operators for the impurity and conduction electron degrees of freedom:
$\hat{d}=\frac{1}{\sqrt{2}}\big(\hat{\eta}_{d}+i\hat{\xi}_{d}\big)\;,\qquad\hat{c}_{k}=\frac{1}{\sqrt{2}}\big(\hat{\eta}_{k}+i\hat{\xi}_{k}\big)\;,$ (242)
where all $\hat{\eta}$ and $\hat{\xi}$ operators are Hermitian Majoranas satisfying $\{\hat{\eta}_{a},\hat{\eta}_{b}\}=\{\hat{\xi}_{a},\hat{\xi}_{b}\}=\delta_{ab}$ and $\{\hat{\eta}_{a},\hat{\xi}_{b}\}=0$. Substituting these definitions into $\hat{H}_{\text{MRLM}}$ yields the full Majorana representation:
$\hat{H}_{\text{MRLM}}=\frac{i\epsilon_{d}}{2}\hat{\eta}_{d}\hat{\xi}_{d}+\frac{i}{2}\sum_{k}\epsilon_{k}\hat{\eta}_{k}\hat{\xi}_{k}+iV\sum_{k}\hat{\eta}_{k}\hat{\xi}_{d}\;.$ (243)
In this basis, the hybridization term couples only one of the impurity Majoranas $\hat{\xi}_{d}$ to the conduction Majoranas $\hat{\eta}_{c}$. This selective coupling captures the physical essence of the 2CK fixed point. Only half of the impurity degree of freedom hybridizes with the conduction bath, while the other half remains free. The MRLM admits a quantum critical point at particle-hole symmetry, $\epsilon_{d}=0$, where the impurity Majorana $\hat{\eta}_{d}$ becomes completely decoupled from the Hamiltonian.
[…]
$T^{*}\sim\frac{\epsilon_{d}^{2}}{\Gamma}\;,$ (244)
where $\Gamma=\pi\rho_{0}V^{2}$ is the hybridization width and $\rho_{0}$ the conduction density of states at the Fermi level. This emergent low-energy scale $T^{*}$ controls the crossover from non-Fermi liquid to Fermi-liquid behavior. For temperatures $T\gg T^{*}$, the system remains in the quantum critical regime characterized by 2CK scaling, whereas for $T\ll T^{*}$ the formerly decoupled Majorana becomes weakly hybridized, restoring full screening and conventional Fermi-liquid behavior.
The energy scale $\Gamma$ plays a role analogous to the Kondo temperature $T_{K}$ setting the energy below which the system flows toward the critical non-Fermi liquid regime. In the low temperature limit $T\ll\Gamma$, all thermodynamic and dynamical quantities exhibit universal scaling governed by this fixed point. Physical observables—such as the impurity spectral function, conductance, or spin susceptibility—show fractional power laws and logarithmic corrections, reflecting the anomalous dimension of the decoupled Majorana and the boundary quantum criticality of the 2CK problem [302].
XI.3 Estimating the on-site potential at the boundary driven critical point
In strongly correlated electronic platforms such as quantum dots, nanowires, and molecular junctions, the natural observables are inherently local and transport based. Nevertheless, the inherent non-Markovian nature of the strong correlations accrued in such systems permits global many-body coherence and entanglement to manifest through such local quantities. In other words, although experimental access is typically restricted to local observables—such as the impurity occupation, current, or conductance—their expectation values and fluctuations retain signatures of the entire correlated environment. In context of critical quantum metrology, we therefore seek to exploit the divergent response of local impurity observables, which faithfully encode the boundary quantum criticality, as opposed to relying explicitly on collective measurements of many-particle states [54, 32, 39].
Figure 19: Estimation of the on-site potential $\epsilon_{d}$ in the 2CK model. (a) QFI $\mathcal{I}_{\epsilon_{d}}$ as a function of on-site potential $\epsilon_{d}$ and temperature $T$ (log-scale). (b) The maximum QFI $\mathcal{I}_{\epsilon_{d}}$ sampled over the temperature window $T\in\{10^{-10},10^{-1}\}$ for a given $\epsilon_{d}$. The critical point occurs at $\epsilon_{d}=0$ and the energy scale is set $\Gamma=1$.
Therefore, we will utilize the impurity quantum dot as an in situ probe to estimate target quantities of interest. In particular, we will consider estimation of the on-site potential $\epsilon_{d}$. Even though the impurity probe constitutes a single local degree of freedom, the reduced density matrix of the dot, denoted as $\hat{\varrho}_{d}$, is strongly influenced by the many-body environment through the hybridization and entanglement with the conduction channels. Crucially, the reduced state of the impurity in the presence of the strongly interacting environment is characterized by a single observable: the average dot occupation $\langle\hat{n}_{d}\rangle=\langle\hat{d}^{\dagger}\hat{d}\rangle$ which determines all elements of $\hat{\varrho}_{d}$ in the basis of impurity states $\{|0\rangle,|1\rangle\}$ as
$\hat{\varrho}_{d}=\big(1-\hat{n}_{d}\big)|0\rangle\langle 0|+\hat{n}_{d}|1\rangle\langle 1|\;.$ (245)
As the system is quadratic, it can be solved exactly by standard Green’s function techniques and this can be exploited to provide a closed form solution of the dot occupation [303]
$\langle\hat{n}_{d}\rangle=\frac{1}{2}-\text{Im}\left\{\frac{\epsilon_{d}}{\pi\Delta}\left[\psi\big(\tfrac{1}{2}+\tfrac{\Gamma+i\Delta}{4\pi T}\big)-\psi\big(\tfrac{1}{2}+\tfrac{\Gamma-i\Delta}{4\pi T}\big)\right]\right\}\;,$ (246)
where $\Delta=\sqrt{4\epsilon_{d}^{2}-\Gamma^{2}}$ and $\psi(\cdot)$ the digamma function. Despite involving only a local expectation value, Eq. (246) encodes the full many-body correlations of the impurity with its reservoirs in the vicinity of the critical point. Moreover, since the impurity is a two-level system whose populations are described completely by the occupation $\langle\hat{n}_{d}\rangle$, the QFI may be expressed as [54, 32]
$\mathcal{I}_{\epsilon_{d}}=\frac{[\partial_{\epsilon_{d}}\langle\hat{n}_{d}\rangle]^{2}}{\langle\hat{n}_{d}\rangle(1-\langle\hat{n}_{d}\rangle)}\;.$ (247)
We plot this quantity in Fig. 19, fixing the energy scale $\Gamma=1$. In panel (a) the metrological sensitivity $\mathcal{I}_{\epsilon_{d}}$ in estimating the on-site potential $\epsilon_{d}$ is shown in the colorscale as a function of on-site potential and temperature. In particular, we note that the QFI diverges at the critical point $\epsilon_{d}=0$ where a metrological enhancement is offered due to boundary criticality, despite making local measurements on the impurity probe $\hat{\varrho}_{d}$. This represents a clear example of critical metrological amplification achieved without collective measurements, arising solely from the divergent response of local observables near criticality—a consequence of diverging correlation length scales which emerge at critical points. For finite $\epsilon_{d}$ the sensitivity becomes reduced. This is because changing the on-site potential induces a change in the emergent energy scale $T^{*}\sim\epsilon_{d}^{2}/\Gamma$—pushing the system out of the critical window.
Panel (b) shows the maximum QFI for each detuning, obtained by sampling over the temperature range in panel (a). This representation highlights the characteristic quantum critical fan—a region of enhanced sensitivity that extends to finite temperature and detuning, bounded by the crossover scale $T^{*}$. Within this fan, the system retains universal non-Fermi liquid correlations and exhibits critical metrological response, while outside this window the behaviour crosses over to conventional Fermi-liquid scaling.
XII Conclusions